Corp.'s breakeven point is revenues of $1,600,000. Fixed costs are $640,000.

Compute the contribution margin percentage.

Compute the selling price if variable costs are ​$15 per unit.

With 70,000 units are sold. Compute the margin of safety in units and dollars.

What does this tell you about the risk of the corp making a​ loss? What are the most likely reasons for this risk to​ increase?
Answer #1

Answer a

At break even point,

Fixed Cost = Contribution

So,

Contribution = $ 640,000

Contribution margin percentage = Contribution / Sales * 100

= $ 640,000 / $1,600,000 = 40%

Answer b

Variable cost is $ 15.

Variable cost ratio = 60% (100 - 40)

Selling price = 15 / 0.6 = $ 25

Answer c

Margin of safety = Sales - Break even sales

Break even units = 64,000 units

Margin of safety (in units) = 70,000 - 64,000 = 6,000 units

Margin of safety (in $) = 6,000 * 25 = $ 150,000

Answer d

The company has a margin cover of $ 150,000.

The most likely reason for this risk to​ increase id decrease in units.

In case of any doubt or clarification, you're welcome to come back via comments.